# Luminance-GS: Adapting 3D Gaussian Splatting to Challenging Lighting Conditions with View-Adaptive Curve Adjustment

### v1

@@ 4.1 Experimental Results
scene metric 3DGS NeRF-W Aleth-NeRF   GS-W Ours “bicycle” PSNR $\uparrow$ 18.87 13.87 11.45 19.00 18.90 SSIM $\uparrow$ 0.618 0.231 0.154 0.603 0.640 LPIPS $\downarrow$ 0.333 0.603 0.772 0.356 0.302 “garden” PSNR $\uparrow$ 19.05 10.71 11.34 20.49 21.01 SSIM $\uparrow$ 0.731 0.222 0.245 0.765 0.786 LPIPS $\downarrow$ 0.242 0.687 0.758 0.222 0.219 “counter” PSNR $\uparrow$ 18.90 14.19 10.11 18.92 19.81 SSIM $\uparrow$ 0.670 0.319 0.187 0.677 0.698 LPIPS $\downarrow$ 0.276 0.641 0.745 0.328 0.271 GPU hrs $\downarrow$ $\sim$ 5min $\sim$ 38hrs $\sim$ 8hrs $\sim$ 45min $\sim$ 14min FPS $\uparrow$ $\sim$ 150 0.138 0.875 $\sim$ 80 $\sim$ 150
@@ 4.2 Ablation Analyze
curve type “buu” (low-light) “buu” (overexposure) “garden” (varying) ${\mathbb{L}}^{g}$ only 17.98/ 0.796/ 0.211 19.00/ 0.789/ 0.343 12.46/ 0.542/ 0.588 ${\mathbb{L}}_{k}^{b}$ only 17.78/ 0.653/ 0.377 16.43/ 0.712/ 0.461 19.05/ 0.777/ 0.256 ${\mathbb{L}}^{g}$ + ${\mathbb{L}}_{k}^{b}$ 18.05/ 0.859/ 0.198 19.37/ 0.804/ 0.344 20.08/ 0.788/ 0.234 ${\mathbb{L}}^{g}$ + ${\mathbb{L}}_{k}^{b}$ + $\mathcal{M}_{k}$ 18.09/ 0.877/ 0.193 19.67/ 0.811/ 0.311 21.01/ 0.786/0.219

### v2

@@ 4.1 Experimental Results
scene metric 3DGS NeRF-W Aleth-NeRF   GS-W Ours “bicycle” PSNR $\uparrow$ 18.63 13.82 11.45 18.33 18.52 SSIM $\uparrow$ 0.539 0.233 0.154 0.618 0.647 LPIPS $\downarrow$ 0.409 0.631 0.772 0.377 0.334 “garden” PSNR $\uparrow$ 20.39 10.71 11.34 20.49 20.93 SSIM $\uparrow$ 0.755 0.222 0.245 0.765 0.786 LPIPS $\downarrow$ 0.205 0.687 0.758 0.222 0.193 “counter” PSNR $\uparrow$ 14.01 10.81 9.87 15.79 16.29 SSIM $\uparrow$ 0.491 0.411 0.249 0.618 0.631 LPIPS $\downarrow$ 0.376 0.758 0.791 0.344 0.313 GPU hrs $\downarrow$ $\sim$ 3min $\sim$ 38hrs $\sim$ 8hrs $\sim$ 45min $\sim$ 18min FPS $\uparrow$ $\sim$ 150 0.138 0.875 $\sim$ 80 $\sim$ 150
@@ 4.2 Ablation Analyze
curve type “buu” (low-light) “buu” (overexposure) “garden” (varying) ${\mathbb{L}}^{g}$ only 17.98/ 0.796/ 0.211 19.00/ 0.789/ 0.343 14.51/ 0.573/ 0.579 ${\mathbb{L}}_{k}^{b}$ only 17.78/ 0.653/ 0.377 16.43/ 0.712/ 0.461 19.15/ 0.729/ 0.251 ${\mathbb{L}}^{g}$ + ${\mathbb{L}}_{k}^{b}$ 18.05/ 0.859/ 0.198 19.37/ 0.804/ 0.344 20.26/ 0.790/ 0.218 ${\mathbb{L}}^{g}$ + ${\mathbb{L}}_{k}^{b}$ + $\mathcal{M}_{k}$ 18.09/ 0.877/ 0.193 19.67/ 0.811/ 0.311 20.93/ 0.786/0.193